OpenStudy (anonymous):

Triangle ABC has side lengths: √6, √2, and 2√2 units. The measures of the angles of the triangle would be... If the base of the triangle has a length of √6 then the measure of the base angle is...

OpenStudy (anonymous):

\[\left(2 \sqrt{2}\right)^2=\sqrt{2}^2+\sqrt{6}^2 \] 8 = 8 So the triangle is a right triangle. Use the Sine Law or trigonometry to calculate one of the acute angles. The other acute angle is 90 degrees or Pi/2 radians minus the first acute angle.